OpenStudy (anonymous):

20! / 16! = A. 24 B. about 16 C. 116,280 im in algebra :)

OpenStudy (anonymous):

c

OpenStudy (anonymous):

How'd u get that answeeeerrrrr? :)

OpenStudy (amistre64):

20.19.18.17

OpenStudy (amistre64):

its what a factorial means 20! means 20.19.18.17.16.16...3.2.1

OpenStudy (anonymous):

\[\frac{20!}{16!}= 20\times 19 \times 18\times 17\] the 16! cancels

OpenStudy (amistre64):

the 16! cancels out the rest from 16 to 1 so we are left with just 20.19.18.17

OpenStudy (anonymous):

Ahhh i see... weird we didnt even learn this in my lesson >.<

OpenStudy (anonymous):

exactly what satellite is expalining

OpenStudy (amistre64):

its secret forbidden knowledge lol

OpenStudy (anonymous):

Well i am probably having a serious *blonde moment* but i have no idea how u got those numberssssss >.<

OpenStudy (anonymous):

ok so the term ! means you multiply all the numbers after that number 20*19*18*17*16 ---------------- 16 The sixteens cancle eachother out so you jsut have to multiply 20*19*18*17=116,280

OpenStudy (anonymous):

Ahh ya and then the 16 cancels out?!?! ohhh i see

OpenStudy (anonymous):

yep